# GraphOracle: A Foundation Model for Knowledge Graph Reasoning

### v1

@@ Appendix F Complete Experimental Results
MRR 0.013 0.032 0.019 0.021 0.075 0.084 0.093 0.425 0.486 0.473 0.493 0.473 0.528 0.539 0.606
H@1 0.028 0.064 0.039 0.042 0.088 0.097 0.104 0.403 0.443 0.428 0.450 0.428 0.486 0.493 0.517
H@10 0.046 0.088 0.048 0.053 0.103 0.125 0.176 0.534 0.566 0.564 0.573 0.586 0.627 0.654 0.756
Table 14: Complete ablation study. $N$ means the number of datasets used in pre-training which can be found in Table 5.
models Nell-100 WK-100 FB-100 YAGO3-10 GeoKG
Type MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
GraphOracle 0.686 0.589 0.898 0.369 0.176 0.693 0.525 0.361 0.799 0.683 0.644 0.798 0.606 0.517 0.756
W/o multi-head 0.605 0.543 0.839 0.317 0.143 0.629 0.498 0.325 0.727 0.639 0.582 0.729 0.574 0.483 0.652
$N=1$ 0.374 0.233 0.541 0.078 0.057 0.125 0.241 0.152 0.389 0.234 0.178 0.356 0.162 0.127 0.285
$N=2$ 0.382 0.211 0.575 0.168 0.123 0.262 0.311 0.190 0.543 0.271 0.193 0.374 0.186 0.154 0.314
$N=3$ 0.397 0.212 0.583 0.172 0.125 0.269 0.321 0.198 0.557 0.297 0.214 0.385 0.192 0.165 0.321
$N=4$ 0.447 0.275 0.624 0.213 0.132 0.314 0.427 0.285 0.593 0.376 0.274 0.413 0.302 0.279 0.436
$N=5$ 0.448 0.277 0.622 0.215 0.136 0.318 0.429 0.286 0.596 0.377 0.275 0.418 0.312 0.288 0.456
$N=6$ 0.450 0.280 0.624 0.215 0.139 0.321 0.432 0.285 0.597 0.375 0.270 0.420 0.311 0.294 0.462
Appendix G Detail Design of GraphOracle+
In real-world scenarios, users often query models with procedural or temporal “How”-type questions rather than isolated factual prompts. Script-based evaluation frameworks [21] emphasize the importance of integrating external knowledge to support dynamic, multi-step reasoning. Motivated by this, we extend GraphOracle by incorporating modality-specific external features to enhance its representational capacity.
[…]
@@ Appendix G Detail Design of GraphOracle+
$\begin{array}[]{ll}&\bm{h}^{0}_{e}(e_{q},r_{q},{c^{i}})=\mathcal{T}(\bm{h}_{e_{i}}(e_{q},r_{q}),c^{i})=\mathcal{T}(\psi(x^{i},c^{i}),c^{i}),\\[5.0pt] &\bm{h}^{\ell}_{e_{o}}(e_{q},r_{q},{c^{i}}^{\ell})=\delta\left(\bm{W}^{\ell}\cdot\sum_{(e_{s},r,e_{o})\in\hat{\mathcal{E}}^{\ell}_{e_{q}}}\alpha^{\ell}_{e_{s},r,e_{o}|r_{q}}\left(\bm{h}^{\ell-1}_{e_{s}}(e_{q},r_{q},{c^{i}}^{\ell-1})+\Psi({\bm{c}^{i}}^{\ell-1},{\bm{c}^{i}}^{\ell},\bm{h}^{\ell}_{r})\right)\right),\par\par\end{array}$ (16)
where ${\alpha}^{\ell}_{e_{s},r,e_{s}|r_{q}}$ is defined the same as Eq. (8), ${\bm{c}^{i}}^{\ell-1}$ and ${\bm{c}^{i}}^{\ell}$ is the modality embedding of nodes at $\ell-1$ and $\ell$ respectively, and $\Psi$ is a vanilla six-layer transformer model for bridging different modalities. Table 1 demonstrates the powerful performance of GraphOracle+ and proves the scalability of our model.
Appendix H Theoretical Analysis of the GraphOracle Model
In this section, we provide rigorous theoretical guarantees for the GraphOracle framework, analyzing its expressiveness, generalization capabilities, convergence properties, stability under perturbations, and relation-dependency correctness.
H.1 Expressiveness and Representation Capacity
@@
$\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\left(\frac{1}{H}\sum_{h=1}^{H}\left[\bm{W}_{1}^{\ell,h}\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\right]\right),$ (17)
@@
Inductive step: Assume the statement holds for L=k. For L=k+1, each relation’s representation now incorporates information from relations that are k+1 steps away. If two relation subgraphs are non-isomorphic within k+1 steps, either: 1. They were already non-isomorphic within k steps, which by our inductive hypothesis leads to different representations, or 2. The difference occurs exactly at step k+1, which will result in different inputs to the (k+1)-th layer message passing function, thus producing different representations.
@@ 1. The RDG representation module:
$\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\left(\frac{1}{H}\sum_{h=1}^{H}\left[\bm{W}_{1}^{\ell,h}\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\right]\right),$ (18)
@@ H.2 Generalization Bounds
$\mathbb{E}[\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]\leq\mathcal{O}\left(\sqrt{\frac{|\mathcal{R}|\cdot d^{2}\cdot L\cdot\log(|\mathcal{R}|\cdot d^{2}\cdot L)}{N}}\right)$ (23)
@@
$\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)=[\mathcal{L}_{test}(\Theta)-\mathcal{L}^{*}_{test}(\Theta)]+[\mathcal{L}^{*}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]$ (26)
@@
H.3 Convergence Analysis
Theorem 5 (Message Passing Convergence).
The RDG message passing in GraphOracle with $L$ layers converges to a unique fixed point as $L\to\infty$ if the maximum eigenvalue of the message propagation matrix $\bm{P}$ is strictly less than 1.
Proof.
We can express the message passing update as:
$\bm{H}^{\ell}=\sigma(\bm{P}\cdot\bm{H}^{\ell-1})$ (27)
where $\bm{H}^{\ell}\in\mathbb{R}^{|\mathcal{R}|\times d}$ is the matrix of all relation representations at layer $\ell$, and $\bm{P}$ is the propagation matrix incorporating the attention weights and transformation matrices.
For convergence, we need to show that this iteration is a contraction mapping. By the Banach fixed-point theorem, if $\bm{P}$ is a contraction, i.e., $\|\bm{P}\|_{2}<1$, then the iteration converges to a unique fixed point.
The attention mechanism ensures that $\sum_{r_{u}\in\mathcal{N}^{past}(r)}\hat{{\alpha}}_{r_{u}r}=1$, which implies that the maximum eigenvalue of the attention matrix is at most 1. The non-linear activation function $\sigma$ is Lipschitz continuous with constant $L_{\sigma}$.
Therefore, the overall iteration is a contraction with constant $L_{\sigma}\cdot\|\bm{P}\|_{2}<1$, guaranteeing convergence to a unique fixed point. ∎
Theorem 6 (Training Convergence Rate).
Under standard assumptions of $L$-smoothness and $\mu$-strong convexity of the loss function, the GraphOracle model trained with gradient descent with learning rate $\eta=\frac{1}{L}$ converges to an $\epsilon$-approximate minimum in $\mathcal{O}(\frac{L}{\mu}\log\frac{1}{\epsilon})$ iterations.
Proof.
Let $\mathcal{L}(\Theta)$ be the loss function, which we assume is $L$-smooth and $\mu$-strongly convex. With gradient descent updates $\Theta_{t+1}=\Theta_{t}-\eta\nabla\mathcal{L}(\Theta_{t})$ and learning rate $\eta=\frac{1}{L}$, we have the following convergence guarantee:
$\mathcal{L}(\Theta_{t})-\mathcal{L}(\Theta^{*})\leq\left(1-\frac{\mu}{L}\right)^{t}\cdot[\mathcal{L}(\Theta_{0})-\mathcal{L}(\Theta^{*})]$ (28)
where $\Theta^{*}$ is the optimal parameter.
To achieve $\mathcal{L}(\Theta_{t})-\mathcal{L}(\Theta^{*})\leq\epsilon$, we need:
$\left(1-\frac{\mu}{L}\right)^{t}\cdot[\mathcal{L}(\Theta_{0})-\mathcal{L}(\Theta^{*})]\leq\epsilon$ (29)
Taking logarithms:
$t\log\left(1-\frac{\mu}{L}\right)+\log[\mathcal{L}(\Theta_{0})-\mathcal{L}(\Theta^{*})]\leq\log\epsilon$ (30)
Using the inequality $\log(1-x)\leq-x$ for $x\in(0,1)$:
$-t\cdot\frac{\mu}{L}+\log[\mathcal{L}(\Theta_{0})-\mathcal{L}(\Theta^{*})]\leq\log\epsilon$ (31)
Solving for $t$:
$t\geq\frac{L}{\mu}\log\frac{\mathcal{L}(\Theta_{0})-\mathcal{L}(\Theta^{*})}{\epsilon}$ (32)
Therefore, the number of iterations required is $\mathcal{O}(\frac{L}{\mu}\log\frac{1}{\epsilon})$. ∎
H.4 Stability and Robustness
Theorem 7 (Lipschitz Continuity).
The GraphOracle model with $L$ layers is $K$-Lipschitz continuous with respect to its inputs, where $K=\prod_{l=1}^{L}\|\bm{W}^{l}\|_{2}\cdot L_{\sigma}^{L}$, and $L_{\sigma}$ is the Lipschitz constant of the activation function.
Proof.
For each layer $\ell$ in the GraphOracle model, the update equation is:
Let $f_{\ell}$ denote the function computed by layer $\ell$. By the chain rule of Lipschitz constants:
$Lip(f_{1}\circ f_{2}\circ\cdots\circ f_{L})\leq\prod_{\ell=1}^{L}Lip(f_{\ell})$ (34)
For each layer $\ell$, we have:
$Lip(f_{\ell})\leq\|\bm{W}^{\ell}\|_{2}\cdot L_{\sigma}$ (35)
where $\|\bm{W}^{\ell}\|_{2}$ is the spectral norm of the weight matrix, and $L_{\sigma}$ is the Lipschitz constant of the activation function.
Therefore, the overall Lipschitz constant of the GraphOracle model is:
$K=\prod_{\ell=1}^{L}\|\bm{W}^{\ell}\|_{2}\cdot L_{\sigma}^{L}$ (36)
This establishes the Lipschitz continuity of the model with respect to its inputs. ∎
Theorem 8 (Perturbation Stability).
Let $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{F})$ and $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{R},\mathcal{F}^{\prime})$ be two KGs differing in at most $k$ facts. Then, for any query $(e_{q},r_{q},?)$:
$\|\bm{h}^{L}(e_{q},r_{q};\mathcal{G})-\bm{h}^{L}(e_{q},r_{q};\mathcal{G}^{\prime})\|_{2}\leq K\cdot\sqrt{k}$ (37)
where $K$ is the Lipschitz constant from the previous theorem.
Proof.
Let $\mathcal{F}\Delta\mathcal{F}^{\prime}$ denote the symmetric difference between the fact sets of the two graphs, with $|\mathcal{F}\Delta\mathcal{F}^{\prime}|\leq k$.
We first analyze how these differences affect the relation graph $\mathcal{G}^{\mathcal{R}}$. Each fact can affect at most $|\mathcal{R}|$ relation edges in $\mathcal{G}^{\mathcal{R}}$. Therefore, the relation graphs $\mathcal{G}^{\mathcal{R}}$ and $\mathcal{G}^{\prime\mathcal{R}}$ differ in at most $k\cdot|\mathcal{R}|$ edges.
For the RDG representation, each differing edge contributes a change of at most $\frac{1}{|\mathcal{N}^{past}(r)|}$ to the representation of a relation. By the Lipschitz continuity of the message passing operations, the total change in relation representations is bounded by:
$\|\bm{R}_{q}-\bm{R}^{\prime}_{q}\|_{F}\leq C_{1}\cdot\sqrt{k\cdot|\mathcal{R}|}$ (38)
where $C_{1}$ is a constant and $\|\cdot\|_{F}$ is the Frobenius norm.
For the universal entity representation, the change in entity representations is proportional to the change in relation representations, bounded by:
$\|\bm{h}^{L}(e_{q},r_{q};\mathcal{G})-\bm{h}^{L}(e_{q},r_{q};\mathcal{G}^{\prime})\|_{2}\leq C_{2}\cdot\|\bm{R}_{q}-\bm{R}^{\prime}_{q}\|_{F}$ (39)
where $C_{2}$ is another constant.
Combining the two bounds and using the Lipschitz constant $K=C_{1}\cdot C_{2}$, we have:
$\|\bm{h}^{L}(e_{q},r_{q};\mathcal{G})-\bm{h}^{L}(e_{q},r_{q};\mathcal{G}^{\prime})\|_{2}\leq K\cdot\sqrt{k}$ (40)
This establishes the stability of the model’s representations under perturbations to the KG. ∎
H.5 Relation-Dependency Correctness
Theorem 9 (Relation-Dependency Path Identification).
Let $r_{i}\rightarrow r_{j}$ denote that relation $r_{i}$ relation-dependencyly influences relation $r_{j}$. If $r_{i}\rightarrow r_{j}$ with influence strength $S(r_{i},r_{j})>\delta$, then as the number of training examples $N\to\infty$, the attention weight $\hat{{\alpha}}_{r_{i}r_{j}}$ in the GraphOracle’s RDG representation satisfies:
$\mathbb{P}(\hat{{\alpha}}_{r_{i}r_{j}}>\frac{\delta}{|\mathcal{N}^{past}(r_{j})|})\to 1$ (41)
Proof.
The multi-head attention weight $\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}$ is computed as:
$\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}=\frac{\exp(\bm{a}^{T}[\bm{W}^{h}_{\alpha}\bm{h}_{r_{u}}^{\ell-1}\|\bm{W}_{\alpha}^{h}\bm{h}_{r_{v}}^{\ell-1}])}{\sum_{r_{w}\in\mathcal{N}^{\text{past}}(r_{v})}\exp(\bm{a}^{T}[\bm{W}_{\alpha}^{h}\bm{h}_{r_{w}}^{\ell-1}\|\bm{W}_{\alpha}^{h}\bm{h}_{r_{v}}^{\ell-1}])},$ (42)
During training, the parameters $a$ and $\bm{W}$ are optimized to minimize the loss function. The relation-dependency influence $r_{i}\rightarrow r_{j}$ with strength $S(r_{i},r_{j})$ means that knowing $r_{i}$ provides information about $r_{j}$ that reduces the uncertainty by $S(r_{i},r_{j})$.
As $N\to\infty$, the trained parameters will reflect this relation-dependency structure, and the model will assign higher attention weights to relation-dependencyly influential relations. Specifically, if $S(r_{i},r_{j})>\delta$, then the model will learn to allocate at least a proportion $\frac{\delta}{|\mathcal{N}^{past}(r_{j})|}$ of the total attention to $r_{i}$.
Therefore, as $N\to\infty$:
$\mathbb{P}(\hat{{\alpha}}_{r_{i}r_{j}}>\frac{\delta}{|\mathcal{N}^{past}(r_{j})|})\to 1$ (43)
This completes the proof. ∎
Theorem 10 (Relation-Dependency Faithfulness).
Let $P(e_{a}|e_{q},r_{q})$ be the true conditional probability of the answer entity given a query, and let $\hat{P}(e_{a}|e_{q},r_{q})$ be the probability estimated by the GraphOracle model. If the model has sufficient capacity and is trained on enough data, then:
$D_{KL}(P(\cdot|e_{q},r_{q})\|\hat{P}(\cdot|e_{q},r_{q}))\leq\epsilon$ (44)
for any query $(e_{q},r_{q},?)$ and any $\epsilon>0$.
Proof.
The GraphOracle model estimates the conditional probability as:
$\hat{P}(e_{a}|e_{q},r_{q})=\frac{\exp(s_{L}(e_{q},r_{q},e_{a}))}{\sum_{e_{o}\in\mathcal{V}}\exp(s_{L}(e_{q},r_{q},e_{o}))}$ (45)
By the universal approximation theorem for neural networks, with sufficient capacity, the model can approximate any continuous function to arbitrary precision. In particular, it can approximate the true scoring function that generates the true conditional probabilities.
Let $s^{*}(e_{q},r_{q},e_{a})$ be the optimal scoring function such that:
$P(e_{a}|e_{q},r_{q})=\frac{\exp(s^{*}(e_{q},r_{q},e_{a}))}{\sum_{e_{o}\in\mathcal{V}}\exp(s^{*}(e_{q},r_{q},e_{o}))}$ (46)
With enough training data and optimization steps, the model’s scoring function $s_{L}(e_{q},r_{q},e_{a})$ will converge to $s^{*}(e_{q},r_{q},e_{a})$ up to an approximation error $\delta$:
$|s_{L}(e_{q},r_{q},e_{a})-s^{*}(e_{q},r_{q},e_{a})|\leq\delta$ (47)
The KL divergence between the true and estimated distributions is bounded by:
$D_{KL}(P(\cdot|e_{q},r_{q})\|\hat{P}(\cdot|e_{q},r_{q}))\leq C\cdot\delta^{2}$ (48)
where $C$ is a constant depending on the structure of the probability distributions.
For any $\epsilon>0$, we can choose $\delta=\sqrt{\frac{\epsilon}{C}}$ and ensure that the model has sufficient capacity to achieve this approximation error, resulting in:
$D_{KL}(P(\cdot|e_{q},r_{q})\|\hat{P}(\cdot|e_{q},r_{q}))\leq\epsilon$ (49)
This establishes the relation-dependency faithfulness of the model. ∎
H.6 Completeness and Structure Preservation
Theorem 11 (Representational Completeness).
For any first-order logic formula $\phi$ over the KG that involves relations $r_{1},r_{2},\ldots,r_{k}$ and entities $e_{1},e_{2},\ldots,e_{m}$, there exists a parameterization of the GraphOracle model that can compute the truth value of $\phi$ with arbitrary precision.
Proof.
We prove this by induction on the structure of first-order logic formulas.
Base case: For atomic formulas $r(e_{i},e_{j})$, which assert that relation $r$ holds between entities $e_{i}$ and $e_{j}$, the GraphOracle model can directly compute its truth value through the score function $s_{L}(e_{i},r,e_{j})$.
Inductive step: Assume that for formulas $\phi_{1}$ and $\phi_{2}$, there exist parameterizations of GraphOracle that compute their truth values with arbitrary precision.
For the logical connectives: 1. Negation: $\neg\phi_{1}$ can be computed as $1-f_{\phi_{1}}$, where $f_{\phi_{1}}$ is the function computing the truth value of $\phi_{1}$. 2. Conjunction: $\phi_{1}\wedge\phi_{2}$ can be computed as $f_{\phi_{1}}\cdot f_{\phi_{2}}$. 3. Disjunction: $\phi_{1}\vee\phi_{2}$ can be computed as $f_{\phi_{1}}+f_{\phi_{2}}-f_{\phi_{1}}\cdot f_{\phi_{2}}$.
For quantifiers: 1. Universal quantification: $\forall x.\phi(x)$ can be approximated by taking the minimum of $f_{\phi}(e)$ over all entities $e$. 2. Existential quantification: $\exists x.\phi(x)$ can be approximated by taking the maximum of $f_{\phi}(e)$ over all entities $e$.
Given that the GraphOracle model can approximate arbitrary continuous functions (as shown in the Expressive Power theorem), it can also approximate these logical operations with arbitrary precision.
Therefore, for any first-order logic formula $\phi$, there exists a parameterization of the GraphOracle model that can compute its truth value with arbitrary precision. ∎
Theorem 12 (Structure-Preserving Transformation).
The transformation from entity-relation interaction space to relation-dependency relation interaction manifold in GraphOracle preserves all path structures up to length $L$, where $L$ is the number of message passing layers.
Proof.
Let $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{F})$ be the original KG and $\mathcal{G}^{\mathcal{R}}=(\mathcal{R},\mathcal{E}^{\mathcal{R}})$ be the transformed relation graph. We need to show that paths in $\mathcal{G}$ correspond to paths in $\mathcal{G}^{\mathcal{R}}$ and vice versa.
First, consider a path $p=(e_{1},r_{1},e_{2},r_{2},\ldots,r_{n-1},e_{n})$ in $\mathcal{G}$ of length $n<L$. This path corresponds to a sequence of relations $r_{1},r_{2},\ldots,r_{n-1}$.
By the definition of $\mathcal{G}^{\mathcal{R}}$, we have:
$(r_{i},r_{i+1})\in\mathcal{E}^{\mathcal{R}}\iff\exists e.(e,r_{i},e^{\prime})\in\mathcal{F}\text{ and }(e^{\prime},r_{i+1},e^{\prime\prime})\in\mathcal{F}$ (50)
Since the path $p$ exists in $\mathcal{G}$, we have $(e_{i},r_{i},e_{i+1})\in\mathcal{F}$ for all $i$. Therefore, $(r_{i},r_{i+1})\in\mathcal{E}^{\mathcal{R}}$ for all $i$, which means that the sequence $r_{1},r_{2},\ldots,r_{n-1}$ forms a path in $\mathcal{G}^{\mathcal{R}}$.
Conversely, consider a path $p^{\prime}=(r_{1},r_{2},\ldots,r_{n})$ in $\mathcal{G}^{\mathcal{R}}$ of length $n<L$. By the definition of $\mathcal{G}^{\mathcal{R}}$, for each $(r_{i},r_{i+1})\in\mathcal{E}^{\mathcal{R}}$, there exist entities $e_{i},e_{i+1},e_{i+2}$ such that $(e_{i},r_{i},e_{i+1})\in\mathcal{F}$ and $(e_{i+1},r_{i+1},e_{i+2})\in\mathcal{F}$.
These facts may not form a single path in $\mathcal{G}$, but they represent possible multi-step reasoning chains. The GraphOracle model can learn to follow these chains through the message passing mechanism in the RDG representation.
Therefore, the transformation preserves all path structures up to length $L$, enabling the model to capture the same reasoning patterns as in the original KG. ∎

### v2

@@ H Detail Design of GraphOracle+
In real-world scenarios, users often query models with procedural or temporal “How”-type questions rather than isolated factual prompts. Script-based evaluation frameworks (li2025scedit) emphasize the importance of integrating external knowledge to support dynamic, multi-step reasoning. Motivated by this, we extend GraphOracle by incorporating modality-specific external features to enhance its representational capacity.
For each entity $e_{i}$, we update its information as $e_{i}=\{x^{i},c^{i}\}$ which combines an additional feature vector $x^{i}$ and a modality tag $c^{i}$. For example, in PrimeKG, a drug may be defined as $c^{i}=\text{"drug"}$ and $x^{i}=\text{"functional description of the drug"}$. To further enhance GraphOracle, we introduce a methodological extension by integrating external information, resulting in an improved variant termed GraphOracle+. Specifically, we incorporate multiple unimodal foundation models (uni-FMs), and by leveraging the embeddings generated by these uni-FMs, we effectively enrich the representations of individual nodes. In the current era of large-scale models, the ability to seamlessly integrate heterogeneous sources of information is of paramount importance. As described, for any two entities $e_{i}$ and $e_{j}$ originating from distinct modalities, we utilize modality-specific foundation models to encode their features. The initial embedding for entity $e_{i}$ under a query context $(e_{q},r_{q})$ is formulated as:
@@
$\begin{array}[]{ll}&\bm{h}^{0}_{e}(e_{q},r_{q},{c^{i}})=\mathcal{T}(\bm{h}_{e_{i}}(e_{q},r_{q}),c^{i})=\mathcal{T}(\psi(x^{i},c^{i}),c^{i}),\\[5.0pt] &\bm{h}^{\ell}_{e_{o}}(e_{q},r_{q},{c^{i}}^{\ell})=\delta\left(\bm{W}^{\ell}\cdot\sum_{(e_{s},r,e_{o})\in\hat{\mathcal{E}}^{\ell}_{e_{q}}}\alpha^{\ell}_{e_{s},r,e_{o}|r_{q}}\left(\bm{h}^{\ell-1}_{e_{s}}(e_{q},r_{q},{c^{i}}^{\ell-1})+\Psi({\bm{c}^{i}}^{\ell-1},{\bm{c}^{i}}^{\ell},\bm{h}^{\ell}_{r})\right)\right),\end{array}$
where ${\alpha}^{\ell}_{e_{s},r,e_{s}|r_{q}}$ is defined the same as Eq. (4.3), ${\bm{c}^{i}}^{\ell-1}$ and $\bm{c}^{i}$$^{\ell}$ is the modality embedding of nodes at $\ell-1$ and $\ell$ respectively, and $\Psi$ is a vanilla six-layer transformer model for bridging different modalities. Table 17 demonstrates the powerful performance of GraphOracle+ and proves the scalability of our model.
I Theoretical Analysis of the GraphOracle Model
In this section, we provide rigorous theoretical guarantees for the GraphOracle framework, analyzing its expressiveness, generalization capabilities, convergence properties, stability under perturbations, and relation‑dependency correctness.
I.1 Expressiveness and Representation Capacity
@@
$\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\left(\frac{1}{H}\sum_{h=1}^{H}\left[\bm{W}_{1}^{\ell,h}\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\right]\right),$
@@
Inductive step: Assume the statement holds for $L=k$. For $L=k+1$, each relation’s representation now incorporates information from relations that are $k+1$ steps away. If two relation subgraphs are non-isomorphic within $k+1$ steps, either:
  1. 1.
    They were already non-isomorphic within $k$ steps, which by our inductive hypothesis leads to different representations, or The difference occurs exactly at step $k+1$, which will result in different inputs to the $(k+1)$-th layer message passing function, thus producing different representations.
@@ 1. The RDG representation module:
$\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\left(\frac{1}{H}\sum_{h=1}^{H}\left[\bm{W}_{1}^{\ell,h}\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\right]\right),$
@@ I.2 Generalization Bounds
$\mathbb{E}[\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]\leq\mathcal{O}\left(\sqrt{\frac{|\mathcal{R}|\cdot d^{2}\cdot L\cdot\log(|\mathcal{R}|\cdot d^{2}\cdot L)}{N}}\right)$
@@
$\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)=[\mathcal{L}_{test}(\Theta)-\mathcal{L}^{*}_{test}(\Theta)]+[\mathcal{L}^{*}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]$
@@
J Comparison of GraphOracle with Supervised SOTA Methods
Dataset # Relation INGRAM ULTRA GraphOracle
NL-25 146 1610 2300 797
NL-50 150 1748 2526 861
NL-75 138 1626 2336 787
NL-100 99 892 1159 416
WK-25 67 598 947 256
WK-50 102 1130 2164 508
WK-75 77 732 1253 313
WK-100 103 1052 1695 460
FB-25 233 7172 10479 3501
FB-50 228 6294 9300 3135
FB-75 213 5042 7375 2524
FB-100 202 4058 5728 2017
WN_V1 9 48 40 37
WN_V2 10 76 76 55
WN_V3 11 94 85 68
WN_V4 9 70 61 54
FB_V1 180 1622 2416 712
FB_V2 200 2692 4050 1237
FB_V3 215 3398 5015 1640
FB_V4 219 4624 7036 2231
NL_V1 14 122 170 51
NL_V2 88 1574 2065 842
NL_V3 142 1942 2558 1017
NL_V4 76 1296 1657 744
(a) KG1
(b) INGRAM
(c) ULTRA
(d) GraphOracle
(e) KG2
(f) INGRAM
(g) ULTRA
(h) GraphOracle
(a) Knowledge Graph
(b) Relation Graph